# PHLP: Sole Persistent Homology for Link Prediction - Interpretable Feature Extraction

### v1

@@ Abstract
Link prediction (LP), inferring the connectivity between nodes, is a significant research area in graph data, where a link represents essential information on relationships between nodes. Although graph neural network (GNN)-based models have achieved high performance in LP, understanding why they perform well is challenging because most comprise complex neural networks. We employ persistent homology (PH), a topological data analysis method that helps analyze the topological information of graphs, to explain the reasons for the high performance. We propose a novel method that employs PH for LP (PHLP) focusing on how the presence or absence of target links influences the overall topology. The PHLP utilizes the angle hop subgraph and new node labeling called degree double radius node labeling (Degree DRNL), distinguishing the information of graphs better than DRNL. Using only a classifier, PHLP performs similarly to state-of-the-art (SOTA) models on most benchmark datasets. Incorporating the outputs calculated using PHLP into the existing GNN-based SOTA models improves performance across all benchmark datasets. To the best of our knowledge, PHLP is the first method of applying PH to LP without GNNs. The proposed approach, employing PH while not relying on neural networks, enables the identification of crucial factors for improving performance.
@@ I Introduction
Figure 2: Topological features in subgraphs with and without a target link $(u,v)$. The diagram illustrates the topological information extraction process for the subgraph $\mathcal{N}$, as described in Section III-D. The presence (top) or absence (bottom) of the target link changes the topological structure of the graph. Top row: When the target link is connected, three features ($C_{1}$, $C_{2}$, and $C_{3}$) are detected shown in the persistence image (PI) in the right column. The PI represents the topological features of the subgraph $\mathcal{N}$ (Section III-E). Bottom row: When the target link is absent, only two features ($C_{2}$ and $C_{3}$) are detected as depicted in the corresponding PI.
In this context, as illustrated in Fig. 1, we present a novel approach to LP, called PHLP, which calculates the topological information of a graph. To use the topological information of subgraphs for LP, we measure how the topological information changes depending on the existence of the target link, as illustrated in Fig. 2. To extract topological information from various perspectives, we utilize angle hop subgraphs for each target node. Additionally, we propose new node labeling called degree double radius node labeling (Degree DRNL), which incorporates degree information for each node, using DRNL [24].
@@
  • We develop an explainable LP method, PHLP, that employs the topological information for LP through PH without relying on neural networks, as illustrated in Fig. 1.
@@ II-A Link Prediction
GNN-Based Methods. The GNN has become a pivotal approach to LP due to its ability to grasp graph-structured data. By effectively incorporating local and global information through message passing and graph aggregation layers, GNNs enhance LP performance. The model by Zhang et el. [24] uses subgraphs as the primary structural units to learn and predict connections, resulting in significant improvement. This paradigm shift led to research focusing on refining and advancing subgraph methods in the context of GNNs [25, 26, 28]. Following this trend, Pan et al. [28] proposed WalkPool (WP), a new pooling mechanism that uses attention to jointly encode node representations and graph topology into learned topological features. However, despite their superior performance, GNN-based methods pose a challenge in comprehending the underlying mechanisms driving their predictions. Within this context, we develop the PHLP, based on PH, with performance comparable to GNN-based models.
@@ II-B Persistent Homology on Graph Data
In recent years, PH, a method of analyzing the topological features of data, has been widely used to analyze graph data. It has demonstrated its effectiveness in graph classification tasks by analyzing the topology of graphs [31, 32, 33, 34, 35, 36, 37, 38] and has been applied to node classification tasks [31, 39, 40]. However, its suitability for LP tasks has been limited, and research on applying PH for LP has progressed slowly. Yan et al. [27] proposed an intriguing approach by integrating PH with GNNs. While their model demonstrates the potential of PH for capturing topological features of graph data, it relies on GNN structures. Additionally, the TLC-GNN requires further research on datasets without node attributes.
@@ III-A Outline of the Proposed Methods
We propose (a) PHLP and (b) multiangle PHLP (MA-PHLP) as described in Fig. 3. The PHLP method analyzes the topological structure of the graph, focusing on target links. First, PHLP samples a $(k,l)$-angle hop subgraph for the given target nodes (Section III-B). Then, PHLP computes persistence images (PIs; Section III-E) for cases with and without the target link. To calculate PIs, we introduce the node labeling and define the edge-weight function (Section III-C). Through PHLP, each target node is transformed into a vector comprising PIs. In addition, LP is performed using the calculated vectors with a classifier (Section III-F). To reflect diverse topological information, we also propose MA-PHLP, which analyzes data from various angles (Section III-G).
@@ III-B Extracting Angle Hop Subgraph
where $d(z,w)$ is the minimum number of edges in any path from $z$ to $w$ in $G$. We define a $(k,l)$-angle hop enclosing subgraph, where the term “angle” signifies viewing the subgraph from multiple perspectives. The (k,l)-angle hop subgraph is a generalization of the $k$-hop subgraph. Given a graph $G=(V,E)$ and two nodes $u,v\in V$, a $(k,l)$-angle hop enclosing subgraph for $(u,v)$ is defined as $\mathcal{N}^{(k,l)}_{u,v}=(V^{\prime},E^{\prime})$ such that
@@
Thus, the angle hop can generate subgraphs in various forms, providing flexibility to adapt to various graph characteristics.
@@ III-C Filtration of the Subgraph
For a given subgraph, the Rips filtration [43, 44, 45] is employed to calculate the topology using PH. To apply the Rips filtration, we define an edge-weight function using node labeling that reflects the topology of the given graph.
@@
Figure 5: Persistence images (PIs) for two node labeling methods for the graphs in Fig. 4. (a) DRNL exhibits identical zero-dimensional PIs for $G_{1}$ and $G_{2}$, (b) Degree DRNL produces distinct outcomes, effectively distinguishing between the two.
@@
However, DRNL encounters limitations when the graph is transformed into node-label information. As depicted in Fig. 4a, DRNL assigns the same node labels to different graphs, resulting in identical zero-dimensional PIs (Fig. 5a, Section III-E). To incorporate the local topology of each node with the effects of DRNL, we introduced Degree DRNL. For a given subgraph $\mathcal{N}=(V^{\prime},E^{\prime})$ of $G$ and center nodes $a,b\in V^{\prime}$, the Degree DRNL $f^{(a,b)}_{\text{degdrnl}}:V^{\prime}\rightarrow\mathbb{R}$ based on $(a,b)$, for all vertex $w$ in $V^{\prime}$, is defined as
$f^{(a,b)}_{\text{degdrnl}}(w)=f^{(a,b)}_{\text{drnl}}(w)+\frac{M-\deg(w)}{M},$
where $M$ denotes the maximum degree of nodes in $\mathcal{N}$. The $(M-\deg(w))/M$ term above assigns larger values for lower degrees of $w$. When $M=\deg(w)$, the value of Degree DRNL matches the original DRNL, ensuring that the edges connected to nodes with higher degrees are assigned smaller values, promoting their earlier emergence in the filtration. Fig. 4b demonstrates various node labels obtained using Degree DRNL, resulting in PIs that can be distinguished from each other (Fig. 5b).
Edge-weight function. For a given subgraph $\mathcal{N}=(V^{\prime},E^{\prime})$, $f:V^{\prime}\rightarrow\mathbb{N}$ denotes any node labeling function. The edge-weight function $W:E^{\prime}\rightarrow\mathbb{R}$, for any edge $(w,z)$ in $E^{\prime}$, is defined as
$W(w,z)=\max(f(w),f(z))+\frac{\min(f(w),f(z))}{\max(f(w),f(z))}.$
The min/max term in the definition of $W$ refines values further, enhancing the discriminative power by reducing the occurrence of identical edge weights.
@@ III-D Persistent Homology
Given an edge-weighted subgraph $\mathcal{N}=(V^{\prime},E^{\prime},W)$, we construct a Rips filtration and compute its PH. First, we create a sequence of subgraphs $\{\mathcal{N}_{\epsilon}\}_{\epsilon\in\mathbb{R}}$, where each $\mathcal{N}_{\epsilon}=(V^{\prime},E^{\prime}_{\epsilon})$ and $E^{\prime}_{\epsilon}=\{e\in E\mid W(e)\leq\epsilon\}$. Second, we convert each subgraph $\mathcal{N}_{\epsilon}$ into the Rips complex $K_{\epsilon}=\{\tau\in\mathbb{X}\mid(w,z)\in E^{\prime}_{\epsilon}\text{ for any two vertices }w,z\in\tau\}$, where $\mathbb{X}$ is the power set of $V^{\prime}$. In $K_{\epsilon}$, a simplex $\tau$ is formed when the vertices in $\tau$ are pairwise connected by edges in $\mathcal{N}_{\epsilon}$. Then, the Rips filtration is obtained as $K_{\epsilon_{1}}\hookrightarrow K_{\epsilon_{2}}\hookrightarrow\cdots\hookrightarrow K_{\epsilon_{m}}=\mathbb{X}$ for $\epsilon_{1}\leq\epsilon_{2}\leq\cdots\leq\epsilon_{m}$. Third, we compute the $p$-dimensional homology group $H_{p}(K_{\epsilon})$ for each complex $K_{\epsilon}$ and track how these groups change as $\epsilon$ increases. The persistence diagram $D$ [45] comprises persistence pairs $(b,d)$ representing the $\epsilon$ values at which a homological feature appears $b$ and disappears $d$, respectively, in the filtration.
III-E Persistence Image
@@
$\rho_{D}(z)=\sum_{(b,d^{\prime})\in L(D)}h(b,d^{\prime})\phi_{(b,d^{\prime})}(z).$
@@
III-F Predicting the Existence of the Target Link
For the given target nodes $(u,v)$, we sample the $(k,l)$-angle hop subgraph $\mathcal{N}^{(k,l)}_{u,v}$, denoted as $\mathcal{N}^{-}$ (Section III-B), assuming that the target link does not exist during this process. On this subgraph, we extract topological features by calculating PH and its vectorization (i.e., the PI, as described in Sections III-D and III-E). The vectorization is calculated for each dimension and concatenated. If $k\neq l$, for symmetry, we repeat the same process with the $(l,k)$-angle hop subgraph once and consider the average of the two vectors, denoting this vector as $x^{-}$. To observe the difference in topological features, we consider a subgraph $\mathcal{N}+$ obtained by connecting the target link to $\mathcal{N}^{-}$. For this graph, $x^{+}$ denotes the vector obtained using this method.

### v2

@@ Abstract
Link prediction (LP), inferring the connectivity between nodes, is a significant research area in graph data, where a link represents essential information on relationships between nodes. Although graph neural network (GNN)-based models have achieved high performance in LP, understanding why they perform well is challenging because most comprise complex neural networks. We employ persistent homology (PH), a topological data analysis method that helps analyze the topological information of graphs, to interpret the features used for prediction. We propose a novel method that employs PH for LP (PHLP) focusing on how the presence or absence of target links influences the overall topology. The PHLP utilizes the angle hop subgraph and new node labeling called degree double radius node labeling (Degree DRNL), distinguishing the information of graphs better than DRNL. Using only a classifier, PHLP performs similarly to state-of-the-art (SOTA) models on most benchmark datasets. Incorporating the outputs calculated using PHLP into the existing GNN-based SOTA models improves performance across all benchmark datasets. To the best of our knowledge, PHLP is the first method of applying PH to LP without GNNs. The proposed approach, employing PH while not relying on neural networks, enables the identification of crucial factors for improving performance.
@@ I Introduction
Figure 2: Topological features in subgraphs with and without a target link $(u,v)$. The diagram illustrates the topological information extraction process for the subgraph $\mathcal{N}$, as described in Section III-D. The presence (top) or absence (bottom) of the target link changes the topological structure of the graph. Top row: When the target link is connected, three features ($C_{1}$, $C_{2}$, and $C_{3}$) are detected shown in the persistence image (PI) in the right column. The PI represents the topological features of the subgraph $\mathcal{N}$ (Section III-F). Bottom row: When the target link is absent, only two features ($C_{2}$ and $C_{3}$) are detected as depicted in the corresponding PI.
In this context, as illustrated in Figure 1, we present a novel approach to LP, called PHLP, which calculates the topological information of a graph. The main difference between GNN-based method and our method is described in Section III-H. To use the topological information of subgraphs for LP, we measure how the topological information changes depending on the existence of the target link, as illustrated in Figure 2. To extract topological information from various perspectives, we utilize angle hop subgraphs for each target node. Additionally, we propose new node labeling called degree double radius node labeling (Degree DRNL), which incorporates degree information for each node, using DRNL [24].
@@
  • We develop an explainable LP method, PHLP, that employs the topological information for LP through PH without relying on neural networks, as illustrated in Figure 1.
@@ II-A Link Prediction
GNN-Based Methods. The GNN has become a pivotal approach to LP due to its ability to grasp graph-structured data. By effectively incorporating local and global information through message passing and graph aggregation layers, GNNs enhance LP performance. The model by Zhang et el. [24] uses subgraphs as the primary structural units to learn and predict connections, resulting in significant improvement. This paradigm shift led to research focusing on refining and advancing subgraph methods in the context of GNNs [25, 26, 28, 43, 44, 45, 46, 47, 48]. However, despite their superior performance, GNN-based methods pose a challenge in comprehending the underlying mechanisms driving their predictions. Within this context, we develop the PHLP, based on PH, with performance comparable to GNN-based models.
@@ II-B Persistent Homology on Graph Data
In recent years, PH, a method of analyzing the topological features of data, has been widely used to analyze graph data. It has demonstrated its effectiveness in graph classification tasks by analyzing the topology of graphs [31, 32, 33, 34, 35, 36, 37, 38] and has been applied to node classification tasks [31, 39, 40]. Bhatia et al. [49] successfully proposed applying PH to LP problem. However, its suitability for LP tasks has been limited, and research on applying PH for LP has progressed slowly. Yan et al. [27] proposed an intriguing approach by integrating PH with GNNs. While their model demonstrates the potential of PH for capturing topological features of graph data, it relies on GNN structures. Additionally, the TLC-GNN requires further research on datasets without node attributes.
@@ III-A Outline of the Proposed Methods
We propose (a) PHLP and (b) multiangle PHLP (MA-PHLP) as described in Figure 3. The PHLP method analyzes the topological structure of the graph, focusing on target links. First, PHLP samples a $(k,l)$-angle hop subgraph for the given target nodes (Section III-B). Then, PHLP computes persistence images (PIs; Section III-F) for cases with and without the target link. To calculate PIs, we introduce the node labeling and define the edge-weight function (Section III-C). Through PHLP, each target node is transformed into a vector comprising PIs. In addition, LP is performed using the calculated vectors with a classifier (Section III-G). To reflect diverse topological information, we also propose MA-PHLP, which analyzes data from various angles (Section III-I).
Figure 4: The motivation of $(k,l)$-angle hop subgraph. Just as viewing photographs of an apple from multiple angles provides a comprehensive understanding. This figure illustrates the capability to extract subgraphs from various perspectives.
@@ III-B Extracting Angle Hop Subgraph
where $d(z,w)$ is the minimum number of edges in any path from $z$ to $w$ in $G$. We define a $(k,l)$-angle hop enclosing subgraph, where the term “angle” signifies viewing the subgraph from multiple perspectives. When defining the $(k,l)$-angle hop subgraph, we were motivated by the observation in Figure 4 that the information captured in a photograph varies depending on the angle from which it is taken. To confirm features like scratches on an apple, photographs from multiple angles are needed for a comprehensive understanding. Similarly, we hypothesized that a multi-perspective approach is necessary to predict the connections between nodes $u$ and $v$ in a graph. Therefore, we devised a method to extract subgraphs by varying the degrees of separation between $u$ and $v$, enabling us to capture different views of the graph.
Given a graph $G=(V,E)$ and two nodes $u,v\in V$, a $(k,l)$-angle hop enclosing subgraph for $(u,v)$ is defined as $\mathcal{N}^{(k,l)}_{u,v}=(V^{\prime},E^{\prime})$ such that
@@
Thus, the angle hop can generate subgraphs in various forms, providing flexibility to adapt to various graph characteristics. The variation in prediction due to angle is discussed in Section IV-C.
@@ III-C Filtration of the Subgraph
For a given subgraph, we construct the filtration to calculate the topology using PH. Existing filtration methods for graphs do not differentiate the target link, which is crucial for predicting the presence of target link. To extract topological features from graphs for this purpose, the filtration process must distinguish the target link. To address this challenge, we utilize a node labeling that marks nodes based on their relative positions to the target nodes, emphasizing the importance of the target link within the graph’s topology. Based on this labeling, we define an edge-weight function that is subsequently used to construct the filtration.
@@
Figure 6: Persistence images (PIs) for two node labeling methods for the graphs in Figure 5. (a) DRNL exhibits identical zero-dimensional PIs for $G_{1}$ and $G_{2}$, (b) Degree DRNL produces distinct outcomes, effectively distinguishing between the two.
@@
However, DRNL encounters limitations when the graph is transformed into node-label information. As depicted in Figure 5a, DRNL assigns the same node labels to different graphs, resulting in identical zero-dimensional PIs (Figure 6a, Section III-F). To incorporate the local topology of each node with the effects of DRNL, we introduced Degree DRNL. For a given subgraph $\mathcal{N}=(V^{\prime},E^{\prime})$ of $G$ and center nodes $a,b\in V^{\prime}$, the Degree DRNL $f^{(a,b)}_{\text{degdrnl}}:V^{\prime}\rightarrow\mathbb{R}$ based on $(a,b)$, for all vertex $w$ in $V^{\prime}$, is defined as
$f^{(a,b)}_{\text{degdrnl}}(w)=f^{(a,b)}_{\text{drnl}}(w)+\frac{M-\deg(w)}{M},$ (2)
where $M$ denotes the maximum degree of nodes in $\mathcal{N}$. The $(M-\deg(w))/M$ term above assigns larger values for lower degrees of $w$. When $M=\deg(w)$, the value of Degree DRNL matches the original DRNL, ensuring that the edges connected to nodes with higher degrees are assigned smaller values, promoting their earlier emergence in the filtration. Figure 5b demonstrates various node labels obtained using Degree DRNL, resulting in PIs that can be distinguished from each other (Figure 6b).
Edge-weight function. For a given subgraph $\mathcal{N}=(V^{\prime},E^{\prime})$, $f:V^{\prime}\rightarrow\mathbb{R}$ denotes any node labeling function. The edge-weight function $W(f):E^{\prime}\rightarrow\mathbb{R}$, for any edge $(w,z)$ in $E^{\prime}$, is defined as $W(f)(w,z)=\max(f(w),f(z)).$
@@ III-D Persistent Homology
Given an edge-weighted subgraph $\mathcal{N}=(V^{\prime},E^{\prime},W)$, we construct a Rips filtration [50, 51, 52] and compute its PH. First, we create a sequence of subgraphs $\{\mathcal{N}_{\epsilon}\}_{\epsilon\in\mathbb{R}}$, where each $\mathcal{N}_{\epsilon}=(V^{\prime},E^{\prime}_{\epsilon})$ and $E^{\prime}_{\epsilon}=\{e\in E\mid W(e)\leq\epsilon\}$. Second, we convert each subgraph $\mathcal{N}_{\epsilon}$ into the Rips complex $K_{\epsilon}=\{\tau\in\mathbb{X}\mid(w,z)\in E^{\prime}_{\epsilon}\text{ for any two vertices }w,z\in\tau\}$, where $\mathbb{X}$ is the power set of $V^{\prime}$. In $K_{\epsilon}$, a simplex $\tau$ is formed when the vertices in $\tau$ are pairwise connected by edges in $\mathcal{N}_{\epsilon}$. Then, the Rips filtration is obtained as $K_{\epsilon_{1}}\hookrightarrow K_{\epsilon_{2}}\hookrightarrow\cdots\hookrightarrow K_{\epsilon_{m}}=\mathbb{X}$ for $\epsilon_{1}\leq\epsilon_{2}\leq\cdots\leq\epsilon_{m}$. Third, we compute the $p$-dimensional homology group $H_{p}(K_{\epsilon})$ for each complex $K_{\epsilon}$ and track how these groups change as $\epsilon$ increases. The persistence diagram $D$ [52] comprises persistence pairs $(b,d)$ representing the $\epsilon$ values at which a homological feature appears $b$ and disappears $d$, respectively, in the filtration.
III-E Theoretical Analysis of the Proposed Filtration Method
We construct filtrations of graphs to compute persistence diagrams and extract topological features. Existing filtrations do not explicitly consider the target link in a graph. We proposed a novel filtration method as explained in Section III-D based on node labeling. The proposed construction demonstrates stability for node labeling functions, providing a strong theoretical foundation for validity.
Definition 1 (Graph isomorphism).
A graph isomorphism between two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a bijection between the vertex sets $f:V_{1}\rightarrow V_{2}$ such that any two vertices $u$ and $v$ of $V_{1}$ are adjacent in $G_{1}$ if and only if $f(u)$ and $f(v)$ are adjacent in $G_{2}$.
When representing graphs as vectors, the goal is to ensure that the same vector represents isomorphic graphs. However, in the context of the LP problem, two isomorphic graphs $G_{1}$ and $G_{2}$ can be represented by different vectors, depending on the placement of the target node. To address this, we redefine graph isomorphism to account for target nodes explicitly. Originally, the target node is needed for calculating the node label, but we denote the target nodes $(u,v)$ in graph $G$ as $G^{(u,v)}$.
Definition 2 (Graph isomorphism with target nodes).
A graph isomorphism with target nodes between two graphs $G_{1}^{(u_{1},v_{1})}$ and $G_{2}^{(u_{2},v_{2})}$ is a graph isomorphism $f$ between $G_{1}$ and $G_{2}$ such that $f(\{u_{1},u_{2}\})=\{v_{1},v_{2}\}$
To guarantee the robustness of our approach under minor perturbations, we present a stability theorem that quantifies the impact of changes in node labeling on the persistence diagrams.
Theorem 1 (Stability theorem).
Let $G^{(u,v)}=(V,E)$ be a graph with target nodes $u,v$ in $V$ and let $f^{(u,v)}$ and $g^{(u,v)}$ be two node labeling fucntions defined on $G^{(u,v)}$. Denote the $p$-dimensional persistence diagrams of $G^{(u,v)}$ obtained from the filtrations constructed by $f^{(u,v)}$ and $g^{(u,v)}$ as $dgm_{p}(f^{(u,v)})$ and $dgm_{p}(g^{(u,v)})$, respectively. Then, the following stability inequality holds:
$D_{B}(dgm_{p}(f^{(u,v)}),dgm_{p}(g^{(u,v)}))\leq\lVert f^{(u,v)}-g^{(u,v)}\rVert_{\infty},$
where $D_{B}$ is bottleneck distance, $\lVert\cdot\rVert_{\infty}$ is infinity norm.
The proof of the theorem is in Appendix A. This theorem aligns with the general framework of stability results for persistence diagrams [53], as established in the context of sub-level set filtrations. Our result extends this framework to node-labeling-based filtrations we designed in the context of graphs, addressing the challenges of incorporating target nodes in LP tasks. It guarantees that small changes in the node labeling function lead to bounded changes in the resulting persistence diagrams, measured by the bottleneck distance. This result validates the robustness of persistence diagrams derived from node labeling functions, ensuring their reliability for LP tasks. In particular, this theorem applies when the node labeling functions are defined as DRNL $f^{(u,v)}_{drnl}$ and Degree DRNL $f^{(u,v)}_{degdrnl}$, as shown in the following corollary.
Corollary 1.
Let $G^{(u,v)}=(V,E)$ be a graph with target nodes $u,v$ in $V$. Then, the bottleneck distance between two persistence diagrams $dgm_{p}(f^{(u,v)}_{drnl})$ and $dgm_{p}(f^{(u,v)}_{degdrnl})$ is bounded by 1.
Proof.
By Theorem 1, we have
$D_{B}(dgm_{p}(f^{(u,v)}_{drnl}),dgm_{p}(f^{(u,v)}_{degdrnl}))$
$\indent\leq\lVert f^{(u,v)}_{drnl}-f^{(u,v)}_{degdrnl}\rVert_{\infty}=\max_{w}(M-\deg(w))/M\leq 1.$
The boundedness established in Corollary 1 demonstrates that DRNL and Degree DRNL yield similar persistence diagrams. This finding confirms that degree information is included while maintaining the information extracted using the DRNL.
Corollary 2.
Suppose there exists a graph isomorphism with target nodes between two graphs $G_{1}^{(u_{1},v_{1})}$ and $G_{2}^{(u_{2},v_{2})}$. Then, the persistence diagrams of $G_{1}^{(u_{1},v_{1})}$ and $G_{2}^{(u_{2},v_{2})}$, obtained from filtrations constructed using the same fixed node labeling function, are identical.
The result guarantees consistency in the representation of isomorphic graphs that preserve target nodes, ensuring they produce identical persistence diagrams under the same node labeling function.
III-F Persistence Image
@@
$\rho_{D}(z)=\sum_{(b,d^{\prime})\in L(D)}h(b,d^{\prime})\phi_{(b,d^{\prime})}(z).$ (3)
@@
III-G Predicting the Existence of the Target Link
For the given target nodes $(u,v)$, we sample the $(k,l)$-angle hop subgraph $\mathcal{N}^{(k,l)}_{u,v}$, denoted as $\mathcal{N}^{-}$ (Section III-B), assuming that the target link does not exist during this process. On this subgraph, we extract topological features by calculating PH and its vectorization (i.e., the PI, as described in Sections III-D and III-F). The vectorization is calculated for each dimension and concatenated. If $k\neq l$, for symmetry, we repeat the same process with the $(l,k)$-angle hop subgraph once and consider the average of the two vectors, denoting this vector as $x^{-}$. To observe the difference in topological features, we consider a subgraph $\mathcal{N}+$ obtained by connecting the target link to $\mathcal{N}^{-}$. For this graph, $x^{+}$ denotes the vector obtained using this method.